What is the equation of parabola whose axis is parallel to y axis and which passes through (4,5,(-2,11)&(-4,21)

Grade:12

2 Answers

Vikas TU
6 years ago
Dear Student,
let eqn of parabola is ax^2 + bx + c = y
  1.  
  2.  
  3.  
Solving we get, a=1/2 , b=-2 and c=5
so eqn of parabola is :
½ x^2 -2x+5=y
or x^2-4x+10=2y [Answer] 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)
Vikas TU
6 years ago
Dear Student,
let eqn of parabola is ax^2 + bx + c = y
(4,5)=>16a+4b+c=5
(-2,11)=>4a-2b+c=11
(-4,21)=>16a-4b+c=21
Solving we get, a=1/2 , b=-2 and c=5
so eqn of parabola is :
½ x^2 -2x+5=y 
or x^2-4x+10=2y [Answer] 
 
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)